# ———————————————————————————————————–

### v2

@@ Appendix E Proof of Theorem 3 and Proposition 2
    That the first addend is 00 has been shown earlier, while for the second one, it is clear using $\dot{l}=\tilde{\rho}^{\rm ss}(\dot{{\cal B}}_{\alpha}({\rm Id}-{\cal B}_{0})^{-1}(\cdot))$.
@@
Let us now study the growth of the derivatives of ~$\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}(\cdot)$ for $0\leq x\leq n\rightarrow+\infty$:
@@
    ~$\begin{split}\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}(\cdot)&=a(u)l(u)(\tilde{\cal B}_{\alpha,u}(\cdot))x(u)+O(\lambda^{x})\\ &=O\left(\frac{u^{2}}{n}+\lambda^{x}\right);\end{split}$
@@
We remark that in the case where the block ~$\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}(\cdot)$ is the first one, due to the action of $\tilde{\rho}^{\rm ss}(u/\sqrt{n})$, the $0^{\rm th}$-order term becomes $O(u^{2}/n+u\lambda^{x}/\sqrt{n})$ and the $1^{\rm st}$-order one becomes $O(u/\sqrt{n})$, while if the block is not the first one, $\lambda^{x}$ decays exponentially fast in $n$ since $x\geq n^{\gamma}$.
@@
    The term $d^{l}(\tilde{\cal B}_{0,u}^{x})/du^{l}$ is a sum over all possible ways of distributing the derivatives among the factors ~$\tilde{\cal B}_{0,u}$.
    Then one can use the spectral decomposition of the ~$\tilde{\cal B}_{0,u}$ which have not been differentiated and glue together differentiated terms using ${\cal R}_{0,u}$ or turning them into products using the projection $l(u)(\cdot)x(u)$ as we did in the previous items. From Eq. (43), one can see that the blocks with a single derivative of a single term ~$\tilde{\cal B}_{0,u}$ bring a growth of the order $n^{2\epsilon^{\prime}}$, the terms with two derivatives (~¨$\ddot{\tilde{{\cal B}}}_{0,u}+2\dot{\tilde{{\cal B}}}_{0,u}\tilde{{\cal R}}_{0,u}\dot{\tilde{{\cal B}}}_{0,u}$) bring a growth of the order $n^{1/2+\epsilon^{\prime}}$ while all other blocks (with at least $3$ elements), cause a growth of the order at most $n$. Notice that the highest growth is attained by making as many groups of three derivatives as possible and this will be used in obtaining the following estimate.
@@
This implies that the first two derivatives spent on a block of the form ~$\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}$ cause a growth equal to $n^{1/2-\epsilon^{\prime}}$, while the other $m-2$ only brings a growth at most equal to the term in Eq. (45). Notice that it is more convenient to spend them in a way that every block of the form ~$\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}$ has two derivatives, since one has that
@@
is of the order of $O(n^{2\epsilon^{\prime}})$ uniformly in $u$ and it is attained by the term where we spend at least two derivatives in each block of the form ~$\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}$ and the last derivative in any of such blocks. One can see that it is not convenient to use any derivative in the final term ~$\tilde{\cal B}_{0,u}^{x}(\mathbf{1})$, since it is better to contrast the decay induced by the other blocks. To conclude, the term in Eq. (42) grows at most as $n^{(2k+3)\epsilon^{\prime}-1/2}=o(n^{2k\epsilon^{\prime}})$. We proved Proposition 2 and notice that Theorem 3 follows considering $\epsilon^{\prime}=0$. ∎
@@ Appendix F Proof of Corollary 3
Let us define the measurable space given by the set $\Omega=\mathbb{N}^{\mathbb{N}}$ together with the $\sigma$-field ${\cal F}$ generated by cylindrical sets; we can consider on $(\Omega,{\cal F})$ the law $\overline{\nu}_{u}$ of $\{N_{\alpha}:\alpha\in{\cal P}\}$ and the law of $\{N_{\alpha}(n):\alpha\in{\cal P}\}$, which, with a slight abuse of notation, we still denote by $\nu_{u,n}$ as well. We know that for every finite set $A$ of patterns, one has
@@
be its block matrix form with respect to the decomposition of $\mathbb{C}^{D}$ into $\mathbb{C}|\chi^{\rm ss}\rangle$ and its orthogonal complement. Then one can prove that $a=b=0$ by using the fact that the Kraus operators are of the form
@@
where $o(n)$ is uniform in $u$ (and $z$ in compact sets). Notice that both $l_{z}(u,n)(\mathbf{1})$ and $\tilde{\rho}^{\rm ss}(u/\sqrt{n})(x_{z}(u))$ converge to $1$ for $n\rightarrow+\infty$; regarding the behaviour of $\lambda_{z}(u,n)$ consider the Taylor expansion up to second order in $u$ around 00:
@@
We can choose $l_{z}(u,n)$ and $x_{z}(u,n)$ such that $\mbox{Tr}(l_{z}(u,n)x_{z}(u,n))\equiv 1$, therefore differentiating $\mbox{Tr}(l_{z}(u,n)\tilde{\mathcal{T}}_{z,u,n}(x_{z}(u,n)))=\lambda_{z}(u,n)$ at 00 one gets
@@
which can be easily seen to be 00. Summing up, we proved that
@@
uniformly in $u$ and in $z$ in compact small neighborhoods of 00. Since for every $u$, $f_{n}(z):=\mathbb{E}_{\nu_{u,n}}[e^{zN_{\rm tot}(n)}]$ are analytic functions around 00, then we can deduce that $e^{\lambda_{z}^{(2)}}$ is analytic as well and that we have uniform convergence (on compact small neighborhoods of 00) of all the derivatives, which consists exactly in the convergence of moments of all orders and we are done. ∎
@@ Appendix G Achievability of the QCRB under additional assumptions
can be upper bounded uniformly in $\tilde{\theta}_{n}$ by a sequence converging to 00. Indeed, notice that for every $a\in\mathbb{R}$
$\begin{split}\mathbb{E}_{\theta}[e^{ia\sqrt{n}(\hat{\theta}_{n}-\theta)}]&=e^{-\frac{a^{2}}{2f_{\theta}}}\mathbb{P}_{\theta}(\theta\in I_{n})\\ &+\int_{\theta\in I_{n}}p_{\theta}(d\tilde{\theta}_{n})(\mathbb{E}_{\theta}[e^{ia\sqrt{n}(\hat{\theta}_{n}-\theta)}|\tilde{\theta}_{n}]-e^{-\frac{a^{2}}{2f_{\theta}}})\\ &+\int_{\theta\notin I_{n}}p_{\theta}(d\tilde{\theta}_{n})\mathbb{E}_{\theta}[e^{ia\sqrt{n}(\hat{\theta}_{n}-\theta)}|\tilde{\theta}_{n}].\\ \end{split}$
@@
$\overline{N}_{n,\theta,\tilde{\theta},\alpha}\sim\text{Poisson}(\lambda_{\alpha}(\tilde{\theta}(\sqrt{n}(\theta-\tilde{\theta})+\tau_{n})^{2}),\quad\alpha\in\mathcal{P}$
@@
    $\alpha^{(i)}$ does not contain more than $n^{\gamma}$ consecutive 00s for every $i=1,\dots,k$;
@@
We denote by $A(n)$ the set of patterns satisfying $1$ and $B(n)$ the set of patterns satisfying $2$, hence $\tilde{G}(n)=A(n)\cap B(n)$. In order to prove the statement, it suffices to show that
@@
where $C(n)$ is the set of pattern counts such that $|\alpha^{(i)}|\leq\eta\log_{2}(n)$ and $\sum_{i=1}^{k}m_{\alpha^{(i)}}\eta\log_{2}(n)+n^{\eta+\gamma}n>n$. The last term amounts to the probability that a Poisson random variable of parameter $\sum_{|\alpha|\leq\eta\log_{2}(n)}\lambda_{\alpha}(\sqrt{n}(\theta-\tilde{\theta})-\tau_{n})^{2}\lesssim n^{3\epsilon}$ is bigger than something that grows as $n/(\eta\log(n)).$ Such a probability goes to 00 uniformly in $\tilde{\theta}$ if $13\epsilon<$ 1 and we are done.

### v3

@@ Appendix E Proof of Theorem 3 and Proposition 2
    That the first addend is 0 has been shown earlier, while for the second one, it is clear using $\dot{l}=\tilde{\rho}^{\rm ss}(\dot{{\cal B}}_{\alpha}({\rm Id}-{\cal B}_{0})^{-1}(\cdot))$.
@@
Let us now study the growth of the derivatives of $\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}(\cdot)$ for $0\leq x\leq n\rightarrow+\infty$:
@@
    $\begin{split}\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}(\cdot)&=a(u)l(u)(\tilde{\cal B}_{\alpha,u}(\cdot))x(u)+O(\lambda^{x})\\ &=O\left(\frac{u^{2}}{n}+\lambda^{x}\right);\end{split}$
@@
We remark that in the case where the block $\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}(\cdot)$ is the first one, due to the action of $\tilde{\rho}^{\rm ss}(u/\sqrt{n})$, the $0^{\rm th}$-order term becomes $O(u^{2}/n+u\lambda^{x}/\sqrt{n})$ and the $1^{\rm st}$-order one becomes $O(u/\sqrt{n})$, while if the block is not the first one, $\lambda^{x}$ decays exponentially fast in $n$ since $x\geq n^{\gamma}$.
@@
    The term $d^{l}(\tilde{\cal B}_{0,u}^{x})/du^{l}$ is a sum over all possible ways of distributing the derivatives among the factors $\tilde{\cal B}_{0,u}$.
    Then one can use the spectral decomposition of the $\tilde{\cal B}_{0,u}$ which have not been differentiated and glue together differentiated terms using ${\cal R}_{0,u}$ or turning them into products using the projection $l(u)(\cdot)x(u)$ as we did in the previous items. From Eq. (45), one can see that the blocks with a single derivative of a single term $\tilde{\cal B}_{0,u}$ bring a growth of the order $n^{2\epsilon^{\prime}}$, the terms with two derivatives ($\ddot{\tilde{{\cal B}}}_{0,u}+2\dot{\tilde{{\cal B}}}_{0,u}\tilde{{\cal R}}_{0,u}\dot{\tilde{{\cal B}}}_{0,u}$) bring a growth of the order $n^{1/2+\epsilon^{\prime}}$ while all other blocks (with at least $3$ elements), cause a growth of the order at most $n$. Notice that the highest growth is attained by making as many groups of three derivatives as possible and this will be used in obtaining the following estimate.
@@
This implies that the first two derivatives spent on a block of the form $\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}$ cause a growth equal to $n^{1/2-\epsilon^{\prime}}$, while the other m2m-2 only brings a growth at most equal to the term in Eq. (47). Notice that it is more convenient to spend them in a way that every block of the form $\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}$ has two derivatives, since one has that
@@
is of the order of $O(n^{2\epsilon^{\prime}})$ uniformly in $u$ and it is attained by the term where we spend at least two derivatives in each block of the form $\tilde{\cal B}_{0,u}^{x}\tilde{\cal B}_{\alpha,u}$ and the last derivative in any of such blocks. One can see that it is not convenient to use any derivative in the final term $\tilde{\cal B}_{0,u}^{x}(\mathbf{1})$, since it is better to contrast the decay induced by the other blocks. To conclude, the term in Eq. (44) grows at most as $n^{(2k+3)\epsilon^{\prime}-1/2}=o(n^{2k\epsilon^{\prime}})$. We proved Proposition 2 and notice that Theorem 3 follows considering $\epsilon^{\prime}=0$. ∎
@@ Appendix F Proof of Corollary 3
[…]
@@
be its block matrix form with respect to the decomposition of $\mathcal{H}_{sa}$ into $\mathbb{C}|\chi^{\rm ss}\rangle$ and its orthogonal complement. Then one can prove that $a=b=0$ by using the fact that the Kraus operators are of the form
@@
where $o(n)$ is uniform in $u$ (and $z$ in compact sets). Notice that both $l_{z}(u,n)(\mathbf{1})$ and $\tilde{\rho}^{\rm ss}(u/\sqrt{n})(x_{z}(u))$ converge to $1$ for $n\rightarrow+\infty$; regarding the behaviour of $\lambda_{z}(u,n)$ consider the Taylor expansion up to second order in $u$ around 0:
@@
We can choose $l_{z}(u,n)$ and $x_{z}(u,n)$ such that $\mbox{Tr}(l_{z}(u,n)x_{z}(u,n))\equiv 1$, therefore differentiating $\mbox{Tr}(l_{z}(u,n)\tilde{\mathcal{T}}_{z,u,n}(x_{z}(u,n)))=\lambda_{z}(u,n)$ at 0 one gets
@@
which can be easily seen to be 0. Summing up, we proved that
@@
uniformly in $u$ and in $z$ in compact small neighborhoods of 0. Since for every $u$, $f_{n}(z):=\mathbb{E}_{\nu_{u,n}}[e^{zN_{\rm tot}(n)}]$ are analytic functions around 0, then we can deduce that $e^{\lambda_{z}^{(2)}}$ is analytic as well and that we have uniform convergence (on compact small neighborhoods of 0) of all the derivatives, which consists exactly in the convergence of moments of all orders and we are done. ∎
@@ Appendix G Achievability of the QCRB under additional assumptions
can be upper bounded uniformly in $\tilde{\theta}_{n}$ by a sequence converging to 0. Indeed, notice that for every $a\in\mathbb{R}$
[…]
@@
$\overline{N}_{n,\theta,\tilde{\theta},\alpha}\sim\text{Poisson}(\lambda_{\alpha}(\tilde{\theta})(\sqrt{n}(\theta-\tilde{\theta})+\tau_{n})^{2}),\quad\alpha\in\mathcal{P}$
@@
    $\alpha^{(i)}$ does not contain more than $n^{\gamma}$ consecutive 0s for every $i=1,\dots,k$;
@@
We denote by $A(n)$ the set of patterns satisfying $1.$ and $B(n)$ the set of patterns satisfying $2.$, hence $\tilde{G}(n)=A(n)\cap B(n)$. In order to prove the statement, it suffices to show that
@@
where $C(n)$ is the set of pattern counts such that $|\alpha^{(i)}|\leq\eta\log_{2}(n)$ and $\sum_{i=1}^{k}m_{\alpha^{(i)}}\eta\log_{2}(n)+n^{\eta+\gamma}n>n$. The last term amounts to the probability that a Poisson random variable of parameter $\sum_{|\alpha|\leq\eta\log_{2}(n)}\lambda_{\alpha}(\sqrt{n}(\theta-\tilde{\theta})-\tau_{n})^{2}\lesssim n^{3\epsilon}$ is bigger than something that grows as $n/(\eta\log(n)).$ Such a probability goes to 0 uniformly in $\tilde{\theta}$ if $3\epsilon<1$ and we are done.